OpenStudy (anonymous):

simplify the numer using the imaginary unit i (square root sign) of -28

OpenStudy (anonymous):

here are the answer choices:

OpenStudy (aakashsudhakar):

√(-28) = √(28) * √(-1) We know that i = √(-1). Therefore, √(28) * i. √(7*4) * i. 2√(7) * i 2i√(7).

OpenStudy (anonymous):

\[a. 2 \sqrt{-7} b. -2\sqrt{-7} c. i \sqrt{28} d. 2i \sqrt{7}\]

OpenStudy (anonymous):

omggg thankyou so much your awesome!!! @AakashSudhakar

OpenStudy (aakashsudhakar):

Anytime!